OpenStudy (anonymous):

Simplify: 36 “a” squared over 7 times 5 over 18 “a”

OpenStudy (anonymous):

36a/(7*5*18) 36a/130

OpenStudy (anonymous):

my answers are A.10a over 7 B 4a^2 over 7 C 10 over 7

OpenStudy (lalaly):

\[\frac{36a^2}{7} \times \frac{5}{18a}=\frac{2a}{7} \times 5 = \frac{10a}{7}\]

OpenStudy (lalaly):

A

OpenStudy (anonymous):

thank you